Find the absolute min and max values of ?(?) = 2?^3 – 32?, on the interval [0, 5].
Answer #1

(x) = 2x²-324 222324 von Ax) = 2x3-32x Differentiase both side worth F(x)= d (213_324) 2 (342) - 3241 fG) = 642-32 for maximf(x) = 22²-32x f(o): 2001 - 3200) & at x = 2.309 FGVE 219.3091?- 32% 2.309. --24.633.6. - 73.000 fox) = -49.2544 & Cet Xrs F(